# Filling the gap in the IERS C01 polar motion series in 1858.9–1860.9

### v2

@@ Abstract
The C01 Earth orientation parameters (EOP) series provided by the International Earth Rotation and Reference Systems Service (IERS) is the longest reliable record of the Earth’s rotation. In particular, the polar motion (PM) series beginning from 1846 provides a basis for investigation of the long-term PM variations. However, the pole coordinate $Y_{p}$ in the IERS C01 PM series has a 2-year gap, which makes this series not completely evenly spaced. This paper presents the results of the first attempt to overcome this problem and discusses some ways to fill this gap. Two novel approaches were considered for this purpose: parametric astronomical model consisting of the bias and the Chandler and annual wobbles with linearly changing amplitudes, and data-driven model based on Singular Spectrum Analysis (SSA). Both methods were tested with various options to ensure robust and reliable results. The results obtained by the two methods generally agree within the $Y_{p}$ errors in the IERS C01 series, but the results obtained by the SSA approach can be considered preferable because it is based on a more complete PM model.
@@ 1 Introduction
Time series of the Earth orientation parameters (EOP) are the main source of information for investigation of the Earth’s rotation, in particular the motion of the Earth’s rotation axis in the terrestrial reference frame, which is also called polar motion (PM). Historical overview of the PM observations and analyses can be found in Dick et al. (2000). The PM is a complex phenomenon, which consists of many variational components including trends and (quasi)periodical oscillations with periods from several hours to several decades. In particular, the study of long-term polar motion variations is crucial for expanding our understanding of Earth’s dynamic processes, particularly in terms of the impacts of climate change, changes in Earth’s internal structure and the redistribution of global mass. To study slow PM variations, long time series of the pole coordinates $X_{p}$ and $Y_{p}$ is needed. The longer the time interval covered by EOP series, the lower frequency variations can be detected and investigated.
@@
The international standard of the EOP series is the combined series computed by the International Earth Rotation and Reference Systems Service (IERS) (Bizouard et al., 2019). The IERS C01 EOP series contains the longest PM series derived from astronomical observations lasting from 1846 to now111ftp://hpiers.obspm.fr/iers/eop/eopc01/eopc01.iau2000.1846-now. It is widely used for analysis of the decadal variations in Earth’s rotation (Gaposchkin, 1972; Vondrák, 1988; Rykhlova and Kurbasova, 1990; Kołaczek and Kosek, 1999; Gambis, 2000; Yatskiv, 2000; Schuh et al., 2001; Höpfner, 2004; Guo et al., 2005; Malkin and Miller, 2010; Vityazev et al., 2010; Zotov, 2010; Miller, 2011; Seitz et al., 2012; Zotov and Bizouard, 2012; Beutler et al., 2020; Lopes et al., 2022; Zotov et al., 2022; Yamaguchi and Furuya, 2024).
@@
However, the IERS PM series has a 2-year gap in $Y_{p}$ pole coordinate in the date range 1858.8 to 1861.0. Of course, a time series without missing epochs is always better than a series with a gap. Missing data can cause problems in statistical data analysis. In particular, for spectral analysis, most of the standard methods used require continuous regularly spaced data. If the latter is the case, the results are free from the risk of introducing spurious frequencies in the spectrum. It is also possible to use other efficient analysis methods for regularly spaced data. The reconstructed continuous evenly spaced PM series can also provide new information about the evolution of the Earth’s rotation modes in the gap epochs.
@@
In this paper, for the first time, to our knowledge, we examine some approaches to address the problem of filling the gap in the IERS C01 series. The challenge is to fill in the missing $Y_{p}$ values in the series while preserving its structure as much as possible. For this purpose, two methods of analysis were considered. The first is the use of a parametric astronomical model, specially developed for this study and consisted of three components: bias, Chandler wobble (CW), and annual wobble (AW) (hereafter referred to as BCA model). The second data-driven model has been constructed using Singular Spectrum Analysis (SSA) (Golyandina et al., 2001). These two methods cover the range of models, from parametric to nonparametric, and allow the treatment of time series that do not have a single fundamental period and are rather short.
@@ 2 Data description and preliminary analysis
In this study, we have focused on the 19th century part of the IERS C01 PM series, primarily on the period 1846–1889. Although the pole coordinates related to the 19th century have relatively low accuracy compared to PM data obtained after 1900, they are of great importance for investigation of long-term variations the Earth’s rotation. Suffice it to say that as early as the late 19th century Chandler (1894a) separated CW and AW signals, derived the value of the CW period of 428-430 days, which is close to the latter determinations, and discovered an $\approx$66-year period of the CW amplitude variation, which is only about 20% shorter than recently obtained from 180-year PM series. Including the 19th century PM data into modern analyses of the entire IERS C01 PM series provides good evidence for  80-year CW periodicity in amplitude and phase variations, which cannot be reliably estimated using only observations since 1900. All this certainly confirms how useful the 19th century PM series are for studying long-term behavior of the PM.
While the work of Fedorov et al. (1972) is widely available and well documented, the original work Rykhlova (1970a) is less known and available and it makes sense to briefly describe it here. This work was based on the analysis of latitude determinations that were made in the framework of several programs for determination of absolute declinations of stars at three observatories: Greenwich, Pulkovo, and Washington. The Greenwich latitude variations were taken from Chandler (1894b) (years 1836.0–1851.0) and Thackeray (1893) (years 1851.0–1891.5). The Pulkovo latitude variations were taken from Orlov (1961) (years 1842.0–1849.5, 1863.6–1879.0, and 1882.0–1892.0). The Washington latitude variations were derived in Rykhlova (1970b) (years 1845.1–1858.8 and 1861.0–1888.0). All three stations were given equal weights.
@@
where $\lambda$ is the latitude of observatory. For most epochs of the Rykhlova PM series, data from all three observatories were used. In the period when Pulkovo was not observing, the pole coordinates were computed using data from Greenwich and Washington. In the Rykhlova series and, consequently, the IERS C01 PM series there are no $Y_{p}$ data from 1858.9 to 1860.9 inclusive. A total of 21 epochs are missing. The middle epoch of the gap is 1859.9. The gap is explained by absence of observations at Washington observatory during this period. The problem is that the Greenwich and Pulkovo observatories are located at close longitudes, $0^{\circ}$ and $30^{\circ}$, respectively, which is close to direction of $X_{p}$. Therefore, it is impossible to accurately estimate the both pole coordinates using only observations from these two stations. Consequently, only $X_{p}$ values were computed for these epochs. The Washington observatory is located at longitude $-77^{\circ}$, which is close to the $Y_{p}$ direction, and its inclusion in the processing provides a good longitude coverage to reliably determine both pole coordinates.
@@ 3 Parametric model
Table 2: Approximation error for the parametric models BCA and BCA2.
@@ 4 SSA analysis
The parameter that we want to optimize using artificial gaps is $r$, the number of the SVD components assigned to the signal. We will call $r$ the model order. To find the optimal value of the model order, we vary it from 1 to 17 and calculate the error of gap filling according to Algorithm 1. We again used RMSE and MAE as a measure of the difference between the imputed values and the input time series values.
@@
Comparison of the data presented in Table 2 and Table 3 shows that the accuracy of the $Y_{p}$ series approximation using CSSA method is a bit better than that for the BCA method. Remarkably, that the best RMSE in filling $\mathsf{Y}$ values achieved by both methods (0.08′′–0.09′′) are very close to the $Y_{p}$ uncertainty given in the IERS C01 series (0.090′′, which is assumed to be the standard deviation, equivalent to RMSE), confirming both the reality of the IERS C01 errors and the correctness of our error estimate. The gap-filling plots with the best parameters described above are shown in Fig. 4.
@@ 6 Conclusions
In this study, mathematical methods were mainly used to complete the existing IERS C01 PM series by filling in the missing $Y_{p}$ values for 21 epochs in the period 1858.9 to 1860.9. On the other hand, mining and reprocessing historical observations looks very interesting, but this huge task goes far beyond the scope of this paper and will obviously require cooperation between observatories that could preserve in their archives declination and latitude observations carried out in the 19th century. Therefore, attempts to improve the quality of the IERS C01 series in the 19th century should be considered, in our opinion, an actual task.

### v3

@@ Abstract
The C01 Earth orientation parameters (EOP) series provided by the International Earth Rotation and Reference Systems Service (IERS) is the longest reliable record of the Earth rotation. In particular, the polar motion (PM) series beginning from 1846 provides a basis for investigation of the long-term PM variations. However, the pole coordinate $Y_{p}$ in the IERS C01 PM series has a 2-year gap, which makes this series not completely evenly spaced. This paper presents the results of the first attempt to overcome this problem and discusses some ways to fill this gap. Two novel approaches were considered for this purpose: parametric astronomical model consisting of the bias and the Chandler and annual wobbles with linearly changing amplitudes, and data-driven model based on Singular Spectrum Analysis (SSA). Both methods were tested with various options to ensure robust and reliable results. The results obtained by the two methods generally agree within the $Y_{p}$ errors in the IERS C01 series, but the results obtained by the SSA approach can be considered preferable because it is based on a more complete PM model.
@@ 1 Introduction
Time series of the Earth orientation parameters (EOP) are the primary source of information for investigation of the Earth rotation, in particular the polar motion (PM). PM observational data are reported by Earth rotation services as the coordinates $X_{p}$ and $Y_{p}$ of the Celestial Intermediate Pole (CIP) in the International Terrestrial Reference System (ITRS). The theoretical foundations of the PM theory are explained, e. g., in Gross (1992); Petit and Luzum (2010) and papers referenced therein. Historical overview of the PM observations and analyses can be found in Dick et al. (2000). The PM is a complex phenomenon, which consists of many variational components including trends and (quasi)periodical oscillations with periods from several hours to several decades. In particular, the study of long-term polar motion variations is crucial for expanding our understanding of Earth’s dynamic processes, particularly in terms of the impacts of climate change, changes in Earth’s internal structure and the redistribution of global mass. To study slow PM variations, long time series of the pole coordinates are needed. The longer the time interval covered by EOP series, the lower frequency variations can be detected and investigated.
@@
The international standard of the EOP series is the combined series computed by the International Earth Rotation and Reference Systems Service (IERS) (Bizouard et al., 2019). The IERS C01 EOP series contains the longest PM series derived from astronomical observations lasting from 1846 to now111ftp://hpiers.obspm.fr/iers/eop/eopc01/eopc01.iau2000.1846-now. It is widely used for analysis of the decadal variations in Earth rotation (Gaposchkin, 1972; Vondrák, 1988; Rykhlova and Kurbasova, 1990; Kołaczek and Kosek, 1999; Gambis, 2000; Yatskiv, 2000; Schuh et al., 2001; Höpfner, 2004; Guo et al., 2005; Malkin and Miller, 2010; Vityazev et al., 2010; Zotov, 2010; Miller, 2011; Seitz et al., 2012; Zotov and Bizouard, 2012; Beutler et al., 2020; Lopes et al., 2022; Zotov et al., 2022; Yamaguchi and Furuya, 2024).
In this study, we have focused on the 19th century part of the IERS C01 PM series, which is very important for analysis of the long-term variations in rotation of the Earth, in particular, Chandler wobble (CW), and annual wobble (AW). Although the pole coordinates related to the 19th century have relatively low accuracy compared to PM data obtained after 1900, they are of great importance for investigation of long-term variations the Earth rotation. Suffice it to say that as early as the late 19th century Chandler (1894a) separated CW and AW signals, derived the value of the CW period of 428-430 days, which is close to the latter determinations, and discovered an $\approx$66-year period of the CW amplitude variation, which is only about 20% shorter than recently obtained from 180-year PM series. Including the 19th century PM data into modern analyses of the entire IERS C01 PM series provides good evidence for  80-year CW periodicity in amplitude and phase variations, which cannot be reliably estimated using only observations since 1900. All this certainly confirms how useful the 19th century PM series is for studying long-term behavior of the PM.
@@
Unfortunately, the IERS C01 series has a 2-year gap in $Y_{p}$ pole coordinate in the date range 1858.9 to 1860.9. Of course, a time series without missing epochs is always better than a series with a gap. Missing data can cause problems in statistical data analysis. In particular, for spectral analysis, most of the standard methods used require continuous regularly spaced data. If the latter is the case, the results are free from the risk of introducing spurious frequencies in the spectrum. It is also possible to use other efficient analysis methods for regularly spaced data. The reconstructed continuous evenly spaced PM series can also provide new information about the evolution of the Earth’s rotation modes in the gap epochs.
@@
In this paper, for the first time, to our knowledge, we examine some approaches to address the problem of filling the gap in the IERS C01 series. The challenge is to fill in the missing $Y_{p}$ values in the series while preserving its structure as much as possible. For this purpose, two methods of analysis were considered. The first is the use of a parametric astronomical model, specially developed for this study and consisted of three components: bias, CW, and AW (hereafter referred to as BCA model). The second data-driven model has been constructed using Singular Spectrum Analysis (SSA) (Golyandina et al., 2001). These two methods cover the range of models, from parametric to nonparametric, and allow the treatment of time series that do not have a single fundamental period and are rather short.
@@ 2 Data description and preliminary analysis
As can be seen from Table 1, the part of the IERS C01 PM series related to 19th century is based on two papers by Rykhlova (1970a) and Fedorov et al. (1972). While the work of Fedorov et al. (1972) is widely available and well documented, the original work Rykhlova (1970a) is less known and available and it makes sense to briefly describe it here. This work was based on the analysis of latitude determinations that were made in the framework of several programs for determination of absolute declinations of stars at three observatories: Greenwich, Pulkovo, and Washington. The Greenwich latitude variations were taken from Chandler (1894b) (years 1836.0–1851.0) and Thackeray (1893) (years 1851.0–1891.5). The Pulkovo latitude variations were taken from Orlov (1961) (years 1842.0–1849.5, 1863.6–1879.0, and 1882.0–1892.0). The Washington latitude variations were derived in Rykhlova (1970b) (years 1845.1–1858.8 and 1861.0–1888.0). All three stations were given equal weights.
@@
where $\Delta\varphi$ is the change of the latitude caused by the PM, and $\lambda$ is the longitude of the observatory. Before computing the pole coordinates, the mean latitude was subtracted from all three latitude series. The same applies to the PM series computed by Fedorov et al. (1972). The mean latitude $\bar{\varphi}$ was computed using the formula applied to latitude time series sampled at 0.1 year:
$\bar{\varphi}_{i}=\left(\sum_{i=0}^{9}\varphi_{i-8}+\varphi_{i-7}+\varphi_{i-2}+\varphi_{i-1}\right)/40\,.$ (2)
As a consequence, the trend component is practically absent in the IERS C01 series before 1900.0. This can be clearly seen in Fig. 1 and is confirmed by further SSA analysis (see Figs. 5 and 6). However, such inconsistency of the two parts of the IERS C01 series before and after 1900.0 with respect to the secular (trend) component does not affect the results of the study of CW and AW variations if an appropriate analysis technique is used. In our case (see the following sections), we use only the data before 1900.0. When using the entire IERS C01 series, the trend component can be removed, for example, using the SSA method or any high-pass filtering.
For most epochs of the Rykhlova’s PM series, data from all three observatories were used. In the period when Pulkovo was not observing, the pole coordinates were computed using data from Greenwich and Washington. In the Rykhlova’s series and, consequently, the IERS C01 PM series there are no $Y_{p}$ data from 1858.9 to 1860.9 inclusive. A total of 21 epochs are missing. The middle epoch of the gap is 1859.9. The gap is explained by absence of observations at Washington observatory during this period. The problem is that the Greenwich and Pulkovo observatories are located at close longitudes, $0^{\circ}$ and $30^{\circ}$, respectively, which is close to direction of $X_{p}$. Therefore, it is impossible to accurately estimate the both pole coordinates using only observations from these two stations. Consequently, only $X_{p}$ values were computed for these epochs. The Washington observatory is located at longitude $-77^{\circ}$, which is close to the $Y_{p}$ direction, and its inclusion in the processing provides a good longitude coverage to reliably determine both pole coordinates.
@@ 3 Parametric model
Table 2: Approximation error for the parametric models BCA and BCA2, arcseconds.
@@ 4 SSA analysis
The parameter that we want to optimize using artificial gaps is $r$, the number of the SVD components assigned to the signal. We will call $r$ the model order. To find the optimal value of the model order, we vary it from 1 to 17 and calculate the error of gap filling according to Algorithm 1 (Appendix 1). We again used RMSE and MAE as a measure of the difference between the imputed values and the input time series values.
@@
Comparison of the data presented in Table 2 and Table 3 shows that the accuracy of the $Y_{p}$ series approximation using CSSA method is a bit better than that for the BCA method. Remarkably, that the best RMSE in filling $\mathsf{Y}$ values achieved by both methods (0.08′′–0.09′′) are very close to the $Y_{p}$ uncertainty given in the IERS C01 series (0.09′′), which is assumed to be the standard deviation, equivalent to RMSE), confirming both the reality of the IERS C01 errors and the correctness of our error estimate. The gap-filling plots with the best parameters described above are shown in Fig. 4.
@@ 5 Final results
As was discussed above, the RMSE of the filling $Y_{p}$ values obtained for both BCA and SSA methods (0.08′′–0.09′′) are very close to the $Y_{p}$ uncertainty in the IERS C01 series (0.09′′). Therefore, it looks reasonable to assign the IERS C01 $Y_{p}$ errors $0.09^{\prime\prime}$ to the filling values given in Table 4.
@@ 6 Conclusions
In this study, mathematical methods were mainly used to complete the existing IERS C01 PM series by filling in the missing $Y_{p}$ values for 21 epochs in the period 1858.9 to 1860.9. Unfortunately, obtained result cannot be verified using the observations employed by Rykhlova (1970a). During the gap period, observations were conducted only by the Greenwich Observatory. However, Greenwich is located at zero longitude, and therefore the Greenwich latitude variations are simply equal to $X_{p}$ (see Eq. (1)), and are thus insensitive to $Y_{p}$.
Generally speaking, attempts to improve the quality of the IERS C01 series in the 19th century, including its extension to the past, should definitely be considered an actual task. In particular, mining and reprocessing of historical observations looks very interesting and important, but this huge task goes far beyond the scope of this paper and will obviously require cooperation between observatories that could preserve in their archives declination and latitude observations carried out in the 19th century. It should also be mentioned that there have been several studies, such as Li et al. (1974); Sekiguchi (1975), attempting to extend the PM series to earlier dates relative to the beginning of the IERS C01 series, but these data do not yet seem reliable enough.
On the other hand, extending the IERS C01 PM series at least 20–25 years into the past, if this is possible at all with acceptable accuracy, is of special interest. The Chandler component of PM shows a minimum amplitude and a phase jump at the epochs around the 1840s, i.e. at the beginning of the IERS C01 series (Malkin and Miller, 2010; Miller, 2011). Therefore, searching for new observational data and refining the pole coordinates for the period 1820–1860 remains a highly desirable goal for improving our knowledge of the features of the CW variations.